OpenStudy (anonymous):

81.5 degrees farenhight equals how much celcious?

OpenStudy (anonymous):

27.5 degrees celcious

OpenStudy (anonymous):

27.5

OpenStudy (anonymous):

convert 8 pounds into appropriate metro units?

OpenStudy (anonymous):

I will tell u One kilogram is equal to: Metric kilotonne 1 *10-06 tonne 0.001 carat 5000 gram 1000 centigram 100000 milligram 1000000 One gram is equal to: Metric kilotonne 1 *10-09 tonne 1 *10-06 centner 1 *10-05 kilogram 0.001 centigram 100 milligram 1000 And also 2.2 pounds is equal to 1 Kg Ur answer 8 Pounds is equal to 1) 3.636 Kg (8/2.2=3.636) 2) 3636g (8/2.2=3.636Kg then 3.636*1000=3636g)

OpenStudy (anonymous):

Convert twenty-one inches to appropriate metric units?

OpenStudy (anonymous):

try it! use the info i just gave you!

OpenStudy (anonymous):

Convert three thousand, eight hundred twenty-four miles to appropriate metric units?

OpenStudy (anonymous):

6,154,131.456 meters.

OpenStudy (anonymous):

Which graph represents the linear relationship that y is six more than negative two times x?

OpenStudy (anonymous):

i cant help with that one..i dont know the options.

OpenStudy (anonymous):

Which of the following is represented by the first coordinates in a group of ordered pairs? The options are: a)relation b)function c)domain d)range

OpenStudy (anonymous):

HELP!!!!!

OpenStudy (anonymous):

domain

OpenStudy (anonymous):

Mark bought several containers of milk at the grocery store for $3 per gallon container. In this situation, what would the number of milk containers be considered? The options are: a)relation b)function c)domain d)range

OpenStudy (anonymous):

HELP!!!!!!!!